# Magnetic Structures Database from Symmetry-aided High-Throughput Calculations

### v2

@@ Abstract
Magnetic structures, which play a central role in determining their physical properties, are known for only very limited compounds. Traditional theoretical methods based on first-principles calculations are fundamentally limited by the need to calculate a large space of input magnetic configurations. Here we introduce a symmetry-aided strategy based on Landau’s phase transition theory. By utilizing the crystallographic space group and the Wyckoff positions of magnetic ions, we narrow down the initial magnetic configurations to a limited number of candidates via the analysis of the group representations. The magnetic ground state is subsequently determined by the lowest energy of those well-seleted magnetic configurations via first-principles calculations. Benchmarking calculations were performed on a subset of the MAGNDATA database with wave vector $\mathbf{q}=0$ and fewer than 40 atoms per unit cell, comprising 302 materials. Our method successfully identified the magnetic structures for 212 of these materials. We further apply our highly efficient method to 7,520 stoichiometric transition metal compounds with fewer than 20 atoms per unit cell in the Inorganic Crystal Structure Database, and establish a magnetic structure database containing 2,900 magnetic materials. To demonstrate the utility of our database, we apply it to the systematic exploration of magnetic topological phases and altermagnets, leading to the identification of 1,070 and 392 instances, respectively.
@@ I Introduction
Magnetic materials are pivotal to modern technology, enabling applications such as information storage, magnetic topology, spintronics and magnonics through their controllable magnetic response Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022). They also occupy a central position in condensed matter physics, providing a fertile platform for exploring fundamental collective phenomena that arise from electronic correlations, symmetry breaking, quantum fluctuations, etc. Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022b, a); Ma et al. (2021). The defining characteristic of a magnetic material is its magnetic structure, the spatial arrangement of magnetic moments, which determines the magnetic space group and strongly influences its physical properties. Establishing the magnetic structure is therefore a prerequisite for studying magnetic phenomena.
@@
While symmetry-based frameworks provide a rigorous and systematic language for describing magnetic configurations Perez-Mato et al. (2024), theoretical prediction of magnetic ground states is most commonly carried out using first-principles total-energy calculations. A fundamental difficulty of such approaches is that they require initial magnetic configurations as input, which in principle have unlimited possibilities. Consequently, the central challenge lies in the absence of an efficient strategy to reduce the vast configuration space to a finite and physically relevant set of candidate magnetic states suitable for first-principles evaluation.
@@
Important progress has been made along this line through a variety of strategies Huebsch et al. (2021); Nomoto et al. (2024); Horton et al. (2019); Zheng and Zhang (2021); Tellez-Mora et al. (2024); Sødequist and Olsen (2024), including constrained searches focusing on collinear magnetism Horton et al. (2019), stochastic optimization schemes such as genetic algorithms that explore magnetic configurations without explicit symmetry assumptions Zheng and Zhang (2021), the linear spin wave theory (LSWT)-based iterative optimization approach Tellez-Mora et al. (2024), and spin-spiral techniques that efficiently capture long-wavelength noncollinear magnetic orderings within simple magnetic unit cells and have successfully determined 194 magnetic structures Sødequist and Olsen (2024). In addition, Huebsch et al. Huebsch et al. (2021); Nomoto et al. (2024) employed a cluster multipole framework and applied it to the AtomWork-Adv database Xu et al. (2011), which contains 228 magnetic compounds. However, magnetic materials exhibit a remarkable diversity of magnetic orderings, including both collinear and noncollinear states. Moreover, the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019) contains on the order of $10^{5}$ materials hosting magnetic ions. These call for a scalable and highly efficient framework capable of treating the full diversity of magnetic orderings, including both collinear and noncollinear states, on an equal footing, irrespective of unit cell complexity.
@@
Within the framework of Landau phase transition theory, we systematically narrow down the vast number of possible initial magnetic configurations to a finite set of candidates based on the irreducible representations (irreps) approach. First-principles total-energy calculations are then employed to identify the energetically favored ground state among these candidate magnetic orders. Benchmark calculations on 302 magnetic structures from MAGNDATA Gallego et al. (2016) show an agreement of about 70%. Applying this approach to compounds with propagation vector $\mathbf{q}=0$ and fewer than 20 atoms per unit cell in ICSD Zagorac et al. (2019), we construct a database comprising over 2,900 magnetic materials. As the magnetic structure fundamentally governs all the physical properties of a material, our magnetic structure database provides a broader materials basis for systematic investigations of physical properties, such as spintronics, magnonics, magnetic topology, multiferroics etc. Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022); Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022b, a); Ma et al. (2021). As an illustration of its capabilities, we present its application to the systematic identification of magnetic topological phases and altermagnets.
@@ II Method
It is well known that most magnetic phase transitions are continuous and can be described within Landau’s theoryAuthier (2014). At the critical temperature, the order parameter emerges continuously from zero, and the resulting magnetic order can be characterized by the irreducible representation of the parent crystallographic space group. This motivates us to construct a set of symmetry-guided magnetic-structure candidates corresponding its irreducible representation. Our approach is as follows.
@@
For a crystal containing $N$ magnetic ions in the unit cell, magnetic configurations with propagation vector $\mathbf{q}=0$ are described by the $3N$ components of the local magnetic moments, $\boldsymbol{\mu}=(m_{1x},m_{1y},m_{1z},\dots,m_{Nx},m_{Ny},m_{Nz})$, where $m_{na}$ denotes the magnetic-moment component of atom $n$ along the crystallographic axis $a$. The $3N$ components of $\boldsymbol{\mu}$ provide a basis for representation $\Gamma_{\mathrm{mag}}$, which is generally reducible and can be decomposed into irreducible representations as
$\Gamma_{\mathrm{mag}}=\sum_{\oplus i}n_{i}\,\Gamma_{i}^{(j)}.$ (1)
@@
Here, $\Gamma_{i}^{(j)}$ denotes an irreducible representation of dimension $j$ from the parent crystallographic space group, and $n_{i}$ is its multiplicity. The completeness relation $\sum_{i}n_{i}\cdot j=3N$ ensures that all magnetic degrees of freedom are fully accounted for in this decomposition.
@@
For a continuous phase transition, the magnetic space groups (MSGs) of the ordered phase must be a subgroup of the parent crystallographic space group. Taking into account the Wyckoff positions of the magnetic atoms, we obtain a specific set of irreducible representations in Eq. (1), thereby reducing the space of candidate MSGs to those compatible with the corresponding instability channels. Namely, a continuous magnetic transition is driven by the instability of a single irreducible representation $\Gamma_{i}^{(j)}$ in Eq. (1), so that only the MSGs associated the irreducible representations appear in Eq. (1) need to be considered. Even after both the unstable irrep and its corresponding MSGs are fixed, the magnetic structure is, in general, still not uniquely determined. To construct explicit magnetic configurations, we therefore employ symmetry-adapted basis vectors of the irreducible representation. However, residual degrees of freedom remain even within this basis-vector description. In the following, we discuss the two distinct sources of this intrinsic non-uniqueness and describe how the associated degrees of freedom can be systematically constrained.
@@
For multi-dimensional irreducible representations $\Gamma_{i}^{(j)}$ with $j>1$, different linear combinations of the basis vectors are symmetry-allowed and may correspond to different MSGs. Although, in principle, the physically realized order parameter could be obtained by minimizing higher-order terms of the Landau free energy, this route is not adopted in the present work. Instead, we employ a symmetry-based criterion and retain, among the magnetic subgroups compatible with the irreducible representations, the one that preserves the highest possible symmetryPerez-Mato et al. (2015); Zhou et al. (2023).
@@
A second source of non-uniqueness arises when a given irreducible representation appears multiple times in the magnetic representation, i.e., $n_{i}>1$ in Eq. (1). In this case, the magnetic structure depends on the particular choice of basis, since multiple symmetry-equivalent basis sets of the basis vectors exist. To eliminate this freedom and construct a finite set of physically realizable configurations, we generate magnetic structures by aligning magnetic moments along all high-symmetry crystallographic directions compatible with the MSGs.
@@
For each candidate material, a finite set of symmetry-allowed magnetic structures associated with the relevant irreducible representations was generated using the procedure described above, thereby drastically reducing the magnetic configuration space to a finite number of initial states for first-principles calculations. First-principles total-energy calculations were then performed among these well-selected structures using the Vienna Ab initio Simulation Package (VASP) Kresse and Joubert (1999); Kresse and Furthmüller (1996), employing a spin-polarized GGA + SOC + $U$ approach. The Coulomb interaction parameters $U$ and $J$ were taken as the averaged values recommended for transition-metal elements based on the linear-response scheme Moore et al. (2024). The ground-state magnetic structure was identified as the configuration with the lowest total energy. To illustrate our approach, the complete workflow with the noncollinear magnet Mn3GaN ,D. and ,E. (1978) is presented in the Supplementary Materials 2. This material is chosen as an example because it exhibits both sources of non-uniqueness discussed above.
We benchmark our framework against established magnetic structures compiled in the MAGNDATA database Gallego et al. (2016). After excluding entries with fractional site occupancies, we restrict our analysis to systems with fewer than 40 atoms in the primitive unit cell and exclude lanthanoids and actinoids due to the challenges in accurately described within standard density functional theory Kotliar et al. (2006). To limit the computational cost at this stage, we focus on magnetic structures with a propagation vector $\mathbf{q}=0$, and this leaves 328 materials for further consideration. After excluding 26 materials whose experimentally reported magnetic structures are incompatible with Landau-type continuous phase transitions (see Supplementary Materials 2), we are left with 302 materials to be validated. By subjecting these materials to our framework, we successfully identify the experimentally reported magnetic ground states of 212 cases, corresponding to an overall accuracy of approximately 70%. Additional details can be found in the Outlook section and the Supplementary Materials 2.
Given that the benchmark yielded excellent results, we applied our framework to the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019). Restricting to systems with unit cells containing fewer than 20 atoms and excluding lanthanoids and actinoids, we focused on $\mathbf{q}=0$ structures and performed systematic high-throughput first-principles calculations to determine the magnetic ground states of 7,520 transition-metal compounds. Within this well-defined scope, our workflow identified 4,619 materials as non-magnetic. The remaining 2,901 magnetic materials were compiled into a magnetic structure database. The magnetic ground states, corresponding magCIF files, and associated information of these magnetic materials have been archived in the Supplementary Materials 2 and a searchable database, available at https://magdatabase.nju.edu.cn.
This database captures a broad diversity of magnetic configurations. Among the identified ground states, 2663 are collinear and 238 are noncollinear. From the viewpoint of magnetic classification, 1,444 materials exhibit ferromagnetic (FM) ordering, and 336 materials are classified as ferrimagnetic (FiM). 1,121 compounds exhibit antiferromagnetic (AFM) order with vanishing net magnetization, including 392 altermagnets. The classification statistics of the magnetic structures are shown in Fig. 1. A complete list of the magnetic materials classification statistics is presented in the Supplementary Materials 2.
@@
Figure 1: Magnetic and topological classification of our magnetic structure database. (a) Magnetic ground-state classification. Among the 2,901 identified magnetic materials, 49.78% exhibit ferromagnetic (FM) ordering, 11.58% are ferrimagnetic (FiM), and the remaining 38.64% are antiferromagnetic (AFM). The FiM materials are further classified into 7.65% collinear (Col) and 3.93% noncollinear (NoC) states according to the relative orientation of magnetic moments. Within the AFM materials, 20.85% exhibit collinear (Col) order, 4.27% exhibit noncollinear (NoC) order, and 13.51% are classified as altermagnets (Alter). (b) Topological classification of magnetic materials. In total, 38% of the magnetic materials in the database are identified as magnetic topological materials, including 4% magnetic topological insulators (TIs) and 34% magnetic topological semimetals (TSMs).
@@
Beyond serving as a catalog of magnetic ground states, we further apply the database to perform topological classification and altermagnets, as will be demonstrated below.

### v3

@@ Abstract
Magnetic structures, which play a central role in determining their physical properties, are known for only very limited compounds. Traditional theoretical approaches to predicting magnetic structures predominantly rely on first-principles calculations. A key challenge of these methods is their requirement for initial magnetic configurations as inputs, which theoretically possess infinite possibilities. In this work, we introduce a strategy based on irreducible representation basis vectors that effectively narrows down the vast space of potential magnetic configurations to a finite set, typically comprising around 20 candidates per material. Despite this significant reduction, the compact input sets generated by our method already encompass the experimental magnetic structures for 253 out of 302 benchmark materials (83.8%) from the MAGNDATA database. These materials have propagation vectors $\mathbf{q}=0$ and unit cells containing up to 40 atoms, all within the Landau framework. Subsequent first-principles calculations correctly identify the magnetic structure in 198 of these cases. We further apply our highly efficient method to 7,520 stoichiometric transition-metal compounds with fewer than 20 atoms per unit cell in the Inorganic Crystal Structure Database, and establish a magnetic structure database containing 2,906 magnetic materials. To demonstrate its utility, we use this database for the systematic exploration of magnetic topological phases and altermagnets, identifying 1,070 and 392 candidate materials, respectively.
@@ I Introduction
Magnetic materials are pivotal to modern technology, enabling applications such as information storage, magnetic topology, spintronics and magnonics through their controllable magnetic response Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022). They also occupy a central position in condensed matter physics, providing a fertile platform for exploring fundamental collective phenomena that arise from electronic correlations, symmetry breaking, quantum fluctuations, etc. Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). The essential characteristic of a magnetic material is its magnetic structure, the spatial arrangement of magnetic moments, which determines the magnetic space group and strongly influences its all physical properties. Establishing the magnetic structure is therefore a prerequisite for studying magnetic phenomena.
@@
While symmetry-based frameworks provide a systematic language for describing magnetic configurations Perez-Mato et al. (2024), theoretical prediction of magnetic structures is most commonly carried out using first-principles total-energy calculations. An essential limitation of such approaches is their dependence on an initial set of magnetic configurations as input, which in principle spans an unbounded space. If the experimental magnetic structure is absent from this initial set, first-principles calculations are fundamentally unable to reproduce it. Consequently, the central challenge lies in the absence of an efficient theoretical strategy to generate a finite set of candidate magnetic configurations that can reliably contain the experimental magnetic structure.
@@
Theoretical progress has been ongoing through a variety of strategies Horton et al. (2019); Zheng and Zhang (2021); Tellez-Mora et al. (2024); Sødequist and Olsen (2024); Huebsch et al. (2021); Nomoto et al. (2024), including constrained searches focusing on collinear magnetism Horton et al. (2019), stochastic optimization schemes such as genetic algorithms that explore magnetic configurations without explicit symmetry assumptions Zheng and Zhang (2021), the linear spin wave theory based iterative optimization approach Tellez-Mora et al. (2024), and spin-spiral techniques that efficiently capture long-wavelength noncollinear magnetic orderings within simple magnetic unit cells and have successfully determined 194 magnetic structures Sødequist and Olsen (2024). In addition, Huebsch et al. Huebsch et al. (2021); Nomoto et al. (2024) employed a cluster multipole framework and applied it to the AtomWork-Adv database Xu et al. (2011), which contains 228 magnetic compounds. However, magnetic materials exhibit a remarkable diversity of magnetic orderings, including both collinear and noncollinear states. Moreover, the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019) contains on the order of $10^{5}$ materials hosting magnetic ions. These call for a scalable and highly efficient framework capable of treating the full diversity of magnetic orderings, including both collinear and noncollinear states, on an equal footing, irrespective of unit cell complexity.
@@
In this work, we develop a magnetic-configuration strategy based on irreducible-representation (irrep) basis vectors. By construction, the resulting initial configuration space is extremely compact, typically comprising only about twenty candidates per material. Despite this drastic reduction, the generated candidate sets already contain the experimental magnetic structures for the majority of benchmark cases. Namely, for 253 out of 302 materials (83.8%) in the MAGNDATA database with propagation vector $\mathbf{q}=0$ and fewer than 40 atoms per unit cell, the experimental magnetic structures are explicitly included. Subsequent first-principles total-energy calculations then correctly identify the magnetic structure in 198 of these cases, corresponding to accuracy of 78.2%. Applying this approach to compounds with propagation vector $\mathbf{q}=0$ and fewer than 20 atoms per unit cell in ICSD Zagorac et al. (2019), we construct a database comprising over 2,900 magnetic materials. As the magnetic structure fundamentally governs all the physical properties of a material, our magnetic structure database provides a broader materials basis for systematic investigations of physical properties, such as spintronics, magnonics, magnetic topology, multiferroics, etc. Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022); Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). As an illustration of its capabilities, we present its application to the systematic identification of magnetic topological phases and altermagnets.
@@ II Method
It is well known that most magnetic phase transitions are continuous and can be described within Landau’s theoryHahn et al. (1983). At the critical temperature, the order parameter emerges continuously from zero, and the resulting magnetic order can be characterized by the irreducible representation of the parent crystallographic space group. This motivates us to construct a set of symmetry-guided magnetic structure candidates corresponding its irreducible representation. Our strategy is as follows.
@@
For a continuous phase transition, the magnetic space groups (MSGs) of the ordered phase must be a subgroup of the parent crystallographic space group. However, the subgroup landscape is typically vast. To systematically reduce the search space of candidate MSGs compared with a brute-force subgroup search, we take the magnetic moments $\boldsymbol{\mu}$ on the magnetic Wyckoff positions as a basis. For a crystal containing $N$ magnetic ions in the unit cell, magnetic configurations with propagation vector $\mathbf{q}=0$ are described by the $3N$ components of the local magnetic moments, $\boldsymbol{\mu}=(m_{1x},m_{1y},m_{1z},\dots,m_{Nx},m_{Ny},m_{Nz})$, where $m_{na}$ denotes the magnetic-moment component of atom $n$ along the crystallographic axis $a$. The $3N$ components of $\boldsymbol{\mu}$ provide a basis for representation $\Gamma_{\mathrm{mag}}$, which is generally reducible and can be decomposed into irreducible representations as
$\Gamma_{\mathrm{mag}}=\sum_{\oplus i}n_{i}\,\Gamma_{i}(j).$ (1)
@@
Here, $\Gamma_{i}(j)$ denotes an irreducible representation of dimension $j$ from the parent crystallographic space group, and $n_{i}$ is its multiplicity. The completeness relation $\sum_{i}n_{i}\cdot j=3N$ ensures that all magnetic degrees of freedom are fully accounted for in this decomposition.
@@
Within Landau theory, a continuous magnetic transition is driven by the instability of a single irreducible representation in Eq. (1). Consequently, only the MSGs compatible with the unstable irrep, namely those that actually appear in Eq. (1), need to be considered. This restriction substantially reduces the search space of candidate MSGs. Even after the unstable irrep and its compatible MSGs are fixed, the magnetic structure is, in general, still not uniquely determined. To construct explicit magnetic configurations, one needs to further specify the symmetry-adapted basis vectors of the irreducible representation. However, even within this basis-vector description, residual degrees of freedom may remain. This non-uniqueness originates from the fact that for multidimensional irreps ($j>1$) and/or repeated irreps ($n_{i}>1$) in Eq. (1), the basis vectors are not uniquely fixed and can be linearly combined in symmetry-allowed ways. In the following, we describe how the associated degrees of freedom can be systematically constrained.
@@
For multi-dimensional irreducible representations $\Gamma_{i}(j)$ with $j>1$ in Eq. (1), the set of the order parameter within the irrep space is not uniquely determined, corresponding to different MSGs. Although, in principle, the physically realized order parameter could be obtained by minimizing higher-order terms of the Landau free energy, this route is not adopted in the present work. Instead, we adopt a symmetry-based criterion and retain all the maximal-symmetry magnetic subgroups among those compatible with the irrep in Eq. (1). Physically, ordering patterns belonging to these MSGs correspond to minimal symmetry breaking at the magnetic transition Perez-Mato et al. (2015); Zhou et al. (2023).
@@
A second source of non-uniqueness arises when a given irreducible representation appears multiple times in the representation, i.e., $n_{i}>1$ in Eq. (1). In this case, the magnetic structure depends on the particular choice of basis, since multiple symmetry-equivalent sets of basis vectors exist and can be linearly combined, giving rise to an infinite number of possible initial input magnetic configurations. To construct a finite set of magnetic configurations, we keep only irreducible-representation basis vectors, whose resulting magnetic moments align with all possible high-symmetry crystallographic directions Hahn et al. (1983) (see Supplementary Materials SM ).
@@
For each candidate material, a finite set of usually $\sim$20 symmetry-allowed magnetic structures associated with the relevant irreducible representations was generated using the procedure described above. First-principles total-energy calculations were then performed among these well-selected magnetic structures using the Vienna Ab initio Simulation Package (VASP) Kresse and Joubert (1999); Kresse and Furthmüller (1996), employing a spin-polarized GGA + SOC + $U$ approach. The Coulomb interaction parameters $U$ and $J$ were taken as the averaged values recommended for transition-metal elements based on the linear-response scheme Moore et al. (2024). The ground-state magnetic structure was identified as the configuration with the lowest total energy.
We benchmark our theory framework against the MAGNDATA database Gallego et al. (2016), focusing on stoichiometric materials with fewer than 40 atoms per unit cell and magnetic propagation vector $\mathbf{q}=0$. Only compounds containing transition-metal ions are considered; lanthanoid and actinoid systems are excluded due to the difficulty of reliably treating within standard density functional theory Kotliar et al. (2006). These criteria yield a total of 328 materials, among which 26 are excluded because their experimental magnetic structures are incompatible with continuous magnetic phase transitions (see Supplementary Materials SM ).
For the remaining 302 materials, our irrep basis-vector strategy systematically reduces the vast space of initial magnetic configurations to a finite set of typically $\sim$20 candidates per material. Remarkably, the experimental magnetic structures of 253 materials (83.8%) are already contained within these candidate sets. Using the generated initial input magnetic configurations ($\sim 5000$ for 253 materials in total), subsequent first-principles total-energy calculations correctly identify the magnetic structures for 198 cases, corresponding to an accuracy of 78.2%.
The noncollinear magnet Mn3GaN Nan et al. (2020) is presented in the Supplementary Materials SM to illustrate our complete workflow. The crystallographic space group of Mn3GaN yields 98 magnetic subgroups and 10 inequivalent irreducible representations. By incorporating the Wyckoff positions of Mn ions, these 10 irreducible representations are reduced to only two relevant ones, $\Gamma_{4}^{+}$ and $\Gamma_{5}^{+}$, together with 9 symmetry-compatible magnetic space groups. Applying the maximal-symmetry criterion further narrows this set to 5 maximal magnetic subgroups. Finally, for the case of repeated irreducible representations ($n_{i}>1$ in Eq. (1)), we retain only basis vectors that generate magnetic moments aligned along high-symmetry crystallographic directions. As a result, the originally vast configuration space is reduced to just 11 candidate magnetic structures, within which the experimental magnetic structure of Mn3GaN is included and subsequently identified by first-principles calculations.
@@
Figure 1: Magnetic and topological classification of our magnetic structure database. (a) Magnetic structure classification. Among the 2,906 identified magnetic materials, 49.86% exhibit ferromagnetic (FM) ordering, 11.63% are ferrimagnetic (FiM), and the remaining 38.51% are antiferromagnetic (AFM). The FiM materials are further classified into 7.67% collinear (Col) and 3.96% noncollinear (NoC) states according to the relative orientation of magnetic moments. Within the AFM materials, 21.54% exhibit collinear (Col) order, 3.48% exhibit noncollinear (NoC) order, and 13.49% are classified as altermagnets (Alter). (b) Topological classification of magnetic materials. In total, 38% of the magnetic materials in the database are identified as magnetic topological materials, including 4% magnetic topological insulators (TIs) and 34% magnetic topological semimetals (TSMs).
Motivated by the excellent benchmark performance, we applied our framework to the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019). Focusing on stoichiometric transition-metal compounds with fewer than 20 atoms per unit cell and excluding lanthanoids and actinoids, we restricted our analysis to magnetic structures with propagation vector $\mathbf{q}=0$. Within this well-defined scope, our irrep basis-vector strategy generates a compact set of trial magnetic configurations, typically <20<20 per material, enabling systematic high-throughput first-principles calculations. Out of 7,520 compounds examined, with a total of approximately 15,000 generated initial magnetic configurations, 4,614 materials are identified as non-magnetic. The remaining 2,906 magnetic materials, together with their magnetic structures, magCIF files, and associated metadata, have been compiled into a searchable database, available at https://magdatabase.nju.edu.cn.
This database contains a broad diversity of magnetic configurations. Among the identified magnetic structures, 2690 are collinear and 216 are noncollinear. From the viewpoint of magnetic classification, 1,449 materials exhibit ferromagnetic (FM) ordering, and 338 materials are classified as ferrimagnetic (FiM). 1,119 compounds exhibit antiferromagnetic (AFM) order with vanishing net magnetization, including 392 altermagnets. The classification statistics of the magnetic structures are shown in Fig. 1. A complete list of the magnetic materials classification statistics is presented in the Supplementary Materials SM .
@@
Beyond serving as a catalog of magnetic structures, we further apply the database to perform topological classification and altermagnets, as will be demonstrated below.